IF a=log 2 3 , b=log 2 5, c=log 7 2 , then find the value of log 140 63 in terms of a,b,c?


IF a=log23 , b=log25, c=log72 , then find the value of  log14063 in terms of a,b,c?


Grade:12th pass

1 Answers

jitender lakhanpal
10 years ago

hi abhishek 

 

log 140 63 = log2 63 / log2 140 =  log2 ( 7 * 32 ) / log2 ( 7 * 5 * 22 )      eqn 1  ---- property of log

and    3  =  2 ^ a           

5 = 2 ^ b    and    7 = 2 ^ (1/c)      ----- given information and definition of logarithm

putting the value in equation 1 

(log2  7 + 3 * log2  3) / (log2 7 + log2  5 +  2 log2 2 ) 

put the values of  3 , 5 , 7 in terms of a , b , c

u will get     {(1/c) + 3a}   /  { (1/c) + b + 2 }